OpenStudy (_.mufasa._):

If a photon has an energy of 8.6 x 10-19 J, what will be its wavelength (in nm)? Show all calculations leading to an answer.

OpenStudy (nechirwan):

The relations are the following: e = h.f (or f = e/h), where: - e is the energy in joules - h is the Planck constant: 6,626 069 572 9.10-34 joule.second - f is the frequency in cycles/second and - \[\lambda\] = c/f , where: - \[\lambda\] = is the wavelength in nanometers - c is the speed of light in vacuum, 299 792 458 m/s And consequently the wavelength is 230,98 nm.